Two kinds of candy sell for 30 pesos and 40 pesos per kilo. How many kilos of each should be used to make 20 kilos of a mixture to sell for 35 pesos per kilo?
Answer by checkley77(12844)

40x+30(20-x)=35*20
40x+600-30x=700
10x=700-600
10x=100
x=100/10
x=10 kilos of 40 pesos is needed.
20-10=10 kilos of 30 pesos is needed.
Proof:
40*10+30*10=35*20
400+300=700
700=700